OpenStudy (anonymous):

Could somebody explain how to use the half-angle identity to find the exact value of sin(5pi/12)?

OpenStudy (p0sitr0n):

So basically you can use the fact that \[\sin(2\alpha)=2\sin(\alpha)\cos(\alpha) \] Then, \[\sin(\frac{5 \pi}{12})\ = 2 \sin (\frac{5 \pi} {6})\cos(\frac{5 \pi}{6})\] But you know the values of both functions at the point \[\alpha = \frac{5 \pi}{6}\] (0.5 and \[\frac{-\sqrt{3}}{2}\]) respectively

OpenStudy (anonymous):

Thank you!